OpenStudy (anonymous):

Suppose 3s represents an even integer. What polynomial represents the product of 3s, the even integer that comes just before 3s, and the even integer that comes just after 3s?

OpenStudy (anonymous):

@jram2326

OpenStudy (anonymous):

Even integers differ by two. If 3s is even, the even just below 3s is (3s -2) and the even just above 3s would be (3s+2). Now, just multiply (3s -2) * (3s) * (3s +2) and see what you get. Start with (3s) * (3s -2) and post what that product is.

OpenStudy (anonymous):

multiply what?

OpenStudy (anonymous):

27s^3 - 12s

OpenStudy (anonymous):

s is any integer. Is s = 3, the 2s = 6. If s = 4, the 2s = 8. The idea is to realize that an even number times an odd = even. Those terms won't go in my calculator except for the calculator on my neck, seriously. If 3s is even, then 3s +2 is even, and the same for 3s -2. Evens differ by a multiple of 2.

OpenStudy (anonymous):

that's what i got too!

OpenStudy (anonymous):

i try let you understand i dont want give you the anwer so fast i want you understand !

OpenStudy (anonymous):

i got 27s^3 - 12s

OpenStudy (anonymous):

Thats the correct answer :)

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

(3s -2) * (3s) * (3s +2) = (3s)(3s - 2) = 9s^2 - 6s (9s^2 -6s) ((3s + 2) = 27s^3 + 18s^2 + ? + ?? (-6s)(3s) = -18s^2 and (-6s)(2) = -12s --------------------- (3s -2) * (3s) * (3s +2) = 27s^3 - 12s---> Answer

OpenStudy (anonymous):

so true midal for you correct!

OpenStudy (anonymous):

thanks @jram2326 you're a life saver! :p

OpenStudy (anonymous):

jram no need give answer we want her to understand !

OpenStudy (anonymous):

Thanks :) Glad to help.

OpenStudy (anonymous):

she had done the problem already, she had the answer. I just confirmed it. But I understand what you're saying.

OpenStudy (anonymous):

he didn't give me answer. I knew it already. I don't really care for math. i hate this subject!

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

no even you are here to understand! we are try to help you even you hate math it eazy to give you all answer without understand we are not here for that!

OpenStudy (anonymous):

I know do you honestly think that I'm that slow!.

OpenStudy (anonymous):

took midal without giving you a detail but it fine nvm

OpenStudy (anonymous):

I don't come on here for answers. I come on here for understanding! and just because I said I hate math, Doesn't mean I come here for answers.! smh

OpenStudy (anonymous):

sorry about that. I'm new to this. I'm don't really know much about the medal thing!

OpenStudy (anonymous):

kk so you should be patient not givin anyone anwser right? i just give you first step to let you do but when you put the anwser i just show you how but nevermind

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

She already had the answer, I dont see why you're making a big deal about me giving it to her.

OpenStudy (anonymous):

jram even she put answer not all people are here to see the answer we are to understand imagine she had wrong answer you plug the correct answer where is the point?

OpenStudy (anonymous):

it an advice i don't mind give her the answer next time let her give the answer then you should if she need help or not :)

OpenStudy (anonymous):

Ok like I said I understand what you are saying. I'm done talking about this.